Question 10 2 pts Questions 10 and 11 are based on the table below, which shows data from a case-control study investigatingInterpret this odds ratio in one sentence.
Answer #1

Odds ratio is the ratio of odds . It refers to the ratio of the odds of an event occurring in the exposed group versus the unexposed group.

Odds ratio = (a/c) / (b/d) or (a×d) / ( b×c)

a = number of exposed to cats in childhood and had allergy = 20

b = number of exposed to cats in childhood, but does not have allergy = 70

c = number of unexposed to cats and have allergy = 80

d= number of unexposed to cats and also does not have allergy = 30

Odds ratio = ( a/c) / ( b/d )

= (20/80 ) / (70/ 30)

= 0.25 / 2.33

0.107

The correct option is 0.11